### v2

# Fully quantum theory of strong-field driven tunable entangled multi-photon states in HHG

Abstract
Quantum high-harmonic generation (HHG) is a prominent and growing field of research with potential capabilities of providing high photon-number entangled states of light. However, there is an open debate regarding the theory level required for correctly describing the quantum aspects of HHG emission, such as squeezing or entanglement. Previous approaches have employed non-interacting classical ensembles of trajectories (sampling the quantum electromagnetic field distribution), or perturbation theory utilizing the classical trajectories as a starting point, missing out key entanglement features. In this Letter, we develop a full quantum theory for entanglement measures in HHG solving exactly the light-matter interaction Hamiltonian (in multiple photonic dimensions) and employ it for evaluating the entanglement between emitted photons of different harmonics. For the first time, we reach qualitative agreement of theory with recent experiments showing that the $R$ entanglement parameter decreases with increasing laser power for below-threshold harmonics [1]. Our results indicate that fine-tuning the laser power could enhance HHG entanglement features, which are observed to oscillate with the driving power and exhibit local non-classical maxima structures. Similarly, our theory predicts that the oscillatory behavior of entanglement observed for below-threshold harmonics also appears for both above-threshold harmonics and between above- and below-threshold harmonics. By analyzing different types of atomic targets, we show that the long-range behavior of driven electronic trajectories can qualitatively change the resulting entanglement, potentially leading to non-universal behavior across systems. Lastly, we show that focal averaging over classical degrees of freedom, which has thus far been ignored in quantum HHG theories, in fact plays a key role in entanglement measures and can change the qualitative behavior of observables such as turning an entangled state into a classical one and vice versa. Our work establishes the state-of-the art in exploring entanglement features in HHG, and paves way for analysis and engineering of ‘truly-quantum’ multi-photon states in the XUV and ultrafast regime for more complex matter systems.
Introduction—High-harmonic generation (HHG) is a nonlinear optical process in which molecules [2, 3], liquids [4, 5], or solids [6] are irradiated by an intense light source and emit higher harmonics of the driving laser. This phenomenon has enabled the birth of new research areas such as attosecond spectroscopy [7, 8], and is routinely used for table-top generation of coherent X-rays [9]. Originally, HHG in atomic and molecular systems was explained through the semiclassical motion of the electron around the nucleus [10], followed by the development a quantum mechanical theory for electronic dynamics [11]. Even though the electronic dynamics were treated using quantum mechanics, the light source and the emitted harmonics were still described with classical electromagnetic theory, preventing the exploration of quantum-optical effects in HHG.
Nevertheless, recent experiments have shown the presence of quantum entanglement in HHG [1, 12] even when matter is irradiated by coherent light. Despite extensive theoretical [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] and experimental [29, 1, 12, 30, 31, 32, 33, 34, 35] advances, the mechanisms that govern entanglement dynamics in strong-field systems remain not fully understood. Previous approaches have employed a range of theories, including analytical perturbative methods [15, 36, 21, 23], as well as semiclassical distribution methods [14, 18, 25, 28] whereby the field is described as a collection of ‘trajectories’ within the framework of classical electromagnetism, and these trajectories are sampled and averaged with proper weights connected to the electromagnetic field’s quasi-probability distribution. While these approximations provide useful insights, their justification is still under investigation. Indeed, we have recently shown that semiclassical approximations for quantum-optical features can lead to very different qualitative HHG spectra for indicent squeezed-coherent light [18]. It is also well-known that semi-classical treatments can miss out key features connected to entanglement in perturbative processes such as spontaneous emission [37] and spontaneous parametric down conversion [38]. In that respect, it would not be surprising that similar issues arise in much higher nonlinear phenomena such as HHG.
Here we develop a comprehensive and formally exact approach that captures entanglement generation between different harmonics in HHG using a single active electron (SAE) driven by coherent light, coupled to multiple quantized harmonic photon modes. Our approach is based on a fully quantum-electrodynamical theory in which the coupling between the electron and the quantized photon modes is solved exactly in higher dimensions using a real-space grid representation, avoiding perturbative or semiclassical approximations [18]. Notably, we find that focal averaging of the laser beam is essential to recover the experimentally observed behavior of the $R$ parameter, which has been neglected thus far in quantum HHG theories. We predict both the magnitude and trend of entanglement between the third and fifth harmonics to be consistent with recent experimental observations in HHG from semiconductors (showing a violation of the Cauchy–Schwarz inequalities [39, 40]), within a similar range of Keldysh parameters [1, 12]. Especially, our theory reproduces qualitative behavior of the $R_{35}$ parameter dropping below the classical limit with increasing laser intensity. Our simulations predict strong oscillations in the $R_{35}$ parameter with the driving laser intensity that can be explained by the modulations of the classical HHG yield, and demonstrate that experimentally tuning the irradiated intensity can be used to maximize entanglement. Similar oscillatory features arise in higher harmonic pairs, suggesting that strong entanglement should be reachable in laser fields with moderate intensities, guiding emerging experiments. Furthermore, by considering different atomic models, we find that long electron trajectories can drastically change entanglement behavior, indicating that the specific treatment of the atomic system plays a key role in quantum correlations in HHG, which might strongly differ between different system (e.g. solid or gas).
Theoretical model and simulation parameters—Our approach consists of a model atom where a single active electron is placed in a soft-Coulomb potential situated at the center of a photonic cavity in the dipole approximation [37, 16, 23, 41, 18]). The overall light-matter Hamiltonian takes the form:
Figure 1: Schematic depiction of the model used to compute the entanglement between two emitted harmonics $\omega_{i}$ and $\omega_{j}$. A gaussian beam of coherent light is radiated to a gas jet of atoms, generating odd harmonics of the incident light. Two integer harmonics of the driving laser frequency $\omega_{L}$ are then selected to perform the photon-counting experiments from which we can recover the instantaneous correlation functions $g^{(2)}_{ij}$ and $R_{ij}$ [see Eq. (3)]
$\begin{split}\hat{H}=\frac{\hat{p}^{2}}{2}-\frac{1}{\sqrt{b^{2}+\hat{x}^{2}}}+\\ +\sum_{n}\frac{4\pi}{L_{C}}\hat{x}^{2}+\sum_{n}\omega_{n}\hat{a}^{\dagger}_{n}\hat{a}_{n}+\hat{H}_{\mathrm{int}}(t),\end{split}$ (1)
$\begin{split}\hat{H}_{\mathrm{int}}(t)=\hat{x}f(t)\left[\sum_{n}\sqrt{\frac{4\pi\omega_{n}}{L_{C}}}\left(\hat{a}_{n}+\hat{a}^{\dagger}_{n}\right)+E_{0}\cos(\omega_{L}t)\right],\end{split}$ (2)
[…]
Optical observables for the photonic sub-system are analyzed at the end of the driving pulse, $t_{f}$, in which the electron has returned to its initial ground-state. This is achieved directly by projecting the combined light-matter system into the electronic ground-state: $|\ket{\phi(t=t_{F})}=\braket{g\|\Psi(t=t_{F})}/|\braket{g\|\Psi(t=t_{F})}$| (see the SI for the explicit projection formula), where the denominator serves for normalization purposes, and is enabled here since the full photonic wave function of these modes is computed exactly. In order to characterize the degree of entanglement in the photonic system, we employ the $R$ parameter [1, 12, 39] that is defined using the creation and annihilation operators of the cavity photon modes $\hat{a}^{(\dagger)}_{i}$:
$R_{ij}=\frac{\braket{\hat{a}^{\dagger}_{i}\hat{a}^{\dagger}_{j}\hat{a}_{i}\hat{a}_{j}}^{2}}{\braket{\hat{a}^{\dagger 2}_{i}\hat{a}^{2}_{i}}\braket{\hat{a}^{\dagger 2}_{j}\hat{a}^{2}_{j}}},$ (3)
where $i$ and $j$ are the harmonic indices. This observable serves as an indicator of entanglement, as its violation of the Cauchy–Schwarz inequalities signals the presence of quantum correlations [1, 39, 40]. Specifically, if $R_{ij}>1$ for two different modes $i$ and $j$, then the system exhibits multipartite entanglement (let us remark that each subsystem super-bunching is not a conclusive criterion for entanglement or squeezing, as we further discuss in the SI). To account for the experimentally spatially varying laser beam profile (where the driving laser interacts with many atomic targets in a gas jet for instance, or with a large volume of a solid), we also perform a focal averaging over the two-dimensional intensity profile of the laser wavefront, modeled by a Gaussian distribution
$I(r)=I_{0}\exp\left[-\frac{r^{2}}{2\sigma^{2}}\right],$ (4)
defined in the radius $0<r<r_{\mathrm{max}}$ for regularization purposes. An illustration of the modeled set-up can be seen in Fig. 1. The intensities considered range from $I_{\mathrm{min}}=I(r_{\mathrm{max}})<I<I_{0}=I(0)$, where $I_{0}$ the peak intensity of the driving beam. Following, the total mixed state of light for intensities given the spatial beam profile is given within that range $I\in(I_{\min},I_{0})$ and is characterized by its peak intensity $I_{0}$:
$\begin{split}\hat{\rho}_{\mathrm{tot}}(I_{0})=\frac{1}{\pi r_{\mathrm{max}}^{2}}\int_{0}^{r_{\mathrm{max}}}dr\int_{0}^{2\pi}rd\theta\hat{\rho}\left[I(r)\right]\\ =\int_{I_{\mathrm{min}}}^{I_{0}}&dIg(I)\hat{\rho}(I),\end{split}$ (5)
$\begin{split}g(I)=\frac{1}{I\ln\left(I_{0}/I_{\mathrm{min}}\right)},\end{split}$ (6)
where $|\hat{\rho}(I)=\ket{\phi(t_{f};I)}\bra{\phi(t_{f};I)}$ is the final photon state driven by an intensity $I$ connected to $E_{0}$. Note that the values of $I<I_{\mathrm{min}}$ are negligible since we expect these observables to vanish as $I\rightarrow 0$. The value of $I_{\mathrm{min}}$ can therefore be decreased until numerical convergence is reached. Using the distribution function from Eq. (6), we evaluate the $R$ parameter for different values of the intensity as:
$\begin{split}R^{\mathrm{av}}_{ij}(I_{0})=\frac{\mathrm{Tr}\left[\hat{\rho}_{\mathrm{tot}}(I_{0})\hat{a}^{\dagger}_{i}\hat{a}^{\dagger}_{j}\hat{a}_{i}\hat{a}_{j}\right]^{2}}{\mathrm{Tr}\left[\hat{\rho}_{\mathrm{tot}}(I_{0})\hat{a}^{\dagger 2}_{i}\hat{a}^{2}_{i}\right]\left[\hat{\rho}_{\mathrm{tot}}(I_{0})\hat{a}^{\dagger 2}_{j}\hat{a}^{2}_{j}\right]}\end{split}$ (7)
where $\hat{\rho}_{\mathrm{tot}}(I_{0})$ is the mixed state with a beam peak power of $I_{0}$ given from Eq. (5). Note that in practice this process ensemble averages the various classical degrees of freedom at the experimental geometry, where a scale of Avogadro number of atoms interact with the beam.
The numerical values of the parameters used in the simulation are driving laser frequency $\omega_{L}=0.057~\mathrm{a.u.}$ (corresponding to $\lambda_{L}=800~\mathrm{nm}$, widely used in HHG [13, 14, 43, 44, 18]), the softening parameter for the electron model potential $b=0.816$ a.u. (corresponding to a Neon ionization potential $I_{p}=0.7925$ a.u.), and the light-matter coupling used is $\lambda=\sqrt{8\pi/L_{C}}=0.01~\mathrm{a.u.}$ (corresponding to a cavity length of $L_{C}\sim 13~\mathrm{\mu m}$, which is consistent with the time-window of the driving pulse, as discussed in the SI). All of the main results and conclusions that will be discussed below arise generally for different values of $\lambda$ (see SI for further results). The minimum intensity for convergence in the focal integration is $I_{\mathrm{min}}=0.32\times 10^{14}~\mathrm{W/cm^{2}}$. All simulations are performed using the Octopus code [45, 46] to solve the Schrödinger equation in a three-dimensional space $(x,y,y^{\prime})$ by expressing photon modes as 1D harmonic oscillators in the photon phase-space (coordinates $y,y^{\prime}$), and the electronic coordinates in real-space ($x$ coordinate) [18]. To suppress unphysical reflections from the boundaries of the simulation box, we apply a complex absorbing potential (CAP), which is converged simultaneously with the box length [47]. In principle, this scheme can be extended into higher dimensions by adding additional photonic coordinates, or electronic coordinates (up to 3D) [48].
Figure 2: Quantum optical observables for third and fifth harmonics in high-harmonic generation (HHG) for different intensities at the end of the driving. Left column displays results for single-intensity simulations $\hat{\rho}(I)$, whereas right column shows focal-averaged results $\hat{\rho}_{\mathrm{tot}}(I_{0})$ vs. laser peak intensity $I_{0}$. (a) Top panels: Expectation values of the four-operator correlators $\braket{\hat{a}^{\dagger}_{i}\hat{a}^{\dagger}_{j}\hat{a}_{i}\hat{a}_{j}}$ for $i=j=3$ (green line), $i=j=5$ (blue line), and $i=3$, $j=5$ (black line). Insets show the corresponding number operator expectation values $\braket{\hat{a}^{\dagger}_{i}\hat{a}_{j}}$. (b) Middle panels: Normalized second-order correlation functions $g^{(2)}_{ij}$. (c) Bottom panels: The $R$ parameter defined in Eq. (3), quantifying non-classical correlations between third and fifth harmonics. Values $R>1$ indicate violation of Cauchy-Schwarz inequalities and signify non-classical correlations.
Figure 3: Focal-averaged $R$ parameter for the 13th and 17th harmonics (top) and the 5th and 17th harmonics (bottom). Values of $R>1$ indicate violation of the Cauchy-Schwarz inequalities, indicating non-classical correlations. The full observable panel for these simulations as the one shown in Fig. 2 is provided in the SI.
Figure 4: Focal-averaged $R$ parameter for four different treatments of the atomic systems: Neon long-range soft-Coulomb potential (blue), Neon soft-Coulomb without long trajectories (green); Neon short-range Gaussian potential, $V(x)=-1.17946\times e^{-0.347x^{2}}$ (red); and a Helium screened soft-Coulomb potential $V(x)=-(1+e^{-x^{2}})/\sqrt{x^{2}+1.9396}$ (cyan). Values of $R>1$ indicate violation of the Cauchy-Schwarz inequalities, indicating non-classical correlations. The full observable panel for this data as the one shown in Fig. 2 is provided in the SI.
Results and discussion—At this stage, we employ the above described numerical approach in order to study the physics and quantum-optical nature of HHG driven by a coherent state of light.
First, we analyze HHG entanglement without focal averaging; i.e., considering just the purely quantum coherent simulation at a single laser intensity $I$, $|\hat{\rho}(I)=\ket{\phi(t_{f};I)}\bra{\phi(t_{f};I)}$, from a single atom. Note that this is the approach that has been employed in all theory works to date, apart from the fact that the quantum degrees of freedom of light were only treated approximately. Our numerical results are shown in the left column of Fig. 2. We find that the values of $R_{35}$ (bottom panel) defined in Eq.(3) between the third and fifth harmonics violate the Cauchy-Schwarz inequalities at certain values of the incident intensity $I$. In addition, we find that $R(I)$ is subject to sharp oscillations with respect to the incident intensity. As shown in the top panel and the inset figure, the expectation values of the mode self-correlations $\braket{\hat{a}^{\dagger 2}_{i}\hat{a}^{2}_{i}}$ and number of photons $\braket{\hat{a}^{\dagger}_{i}\hat{a}_{i}}$ for the emitted harmonics also exhibit oscillations with the driving intensity. These can be interpreted based on channel-closing effects, which have been reported in both theoretical and experimental studies on gases [11, 49, 50, 51] when light is considered classical. These oscillations are also observed in the normalized second-order correlations $g^{(2)}_{ij}$ of the middle panel. Essentially, the oscillations in the emitted photon number (first-order correlations) are also present in the second-order correlations, showing a close connection between the correlation of the emitted photons and the channel-closing effects observed in the dipole yield (which match those of the emitted photon number).
Notably, the extremely sharp oscillations of $R(I)$ in Fig. 2 seem somewhat unphysical, as they suggest an instability or at least a lack of robustness, given that minute changes in laser intensity would strongly affect the outcome. We also note that the intensity dependence predicted here differs significantly from the results reported in quantum HHG for solids [1, 12], raising further doubts regarding their applicability. However, after thorough investigations, we confirmed that this behavior is in fact correct and exact for the studied Hamiltonian of Eq. (1). We suspect that this feature arises since the Hamiltonian is highly quantum-coherent by construction, and observables are therefore sensitive to minute quantum interference effects in harmonic yields, phases, polarizations, etc. In realistic experiments however, additional classical degrees of freedom (most importantly the spatial and temporal laser beam profile) should also play some role, and physically lead to ensemble averaging over an order of Avogadro number of atoms. Such statistical averaging over essentially classical degrees of freedom has not been considered to date in quantum theories of HHG, and we are motivated to test if including it in the theory might suppress such spurious features, and perhaps lead to the correct experimentally observed trend, which is that the $R(I)$ parameter should reduce as the laser intensity increases [1].
We now test the first and second order correlations, as well as the $R$ parameter, with focal averaging (right column of Fig. 2) over a range of intensities as described in Eq. (5). Our results show a completely qualitatively different behavior compared to just the peak laser driving point (left column of Fig. 2). The x-axis of the right column corresponds now to the peak intensity of the laser beam $I_{0}$ from Eq. 6. The focal-averaged $R$ parameter $R^{\mathrm{av}}(I_{0})$ [Eq.(7)] of the bottom panel now exhibits a clear and smoother decreasing trend with the laser peak intensity. The normalized correlation functions $g^{(2)}_{ij}$ in the middle panel show also a decreasing trend with the peak intensity. Both results are consistent with previous experimental results on semiconductors [1, 12] and align with the expectation that increasing laser intensity reduces the bunching of the emitted harmonics, and generally converges to a classical limit. In this case, this reduction of the correlations correspond to an incoherent emission from different atoms. Values of $R>1$, indicating a violation of the Cauchy–Schwarz inequalities [1, 12, 39] and hence the presence of nonclassical correlations (quantum entanglement), are observed within the intensity range $0.7\times 10^{14}~\mathrm{W/cm^{2}}$ to $1.5\times 10^{14}~\mathrm{W/cm^{2}}$ (associated with a Keldysh parameter of 1.01.51.0-1.5). The entanglement reaches a maximum of $1.7$, which is of the same order of magnitude as reported in recent experimental studies [1, 12]. These results demonstrate that truly non-classical light can emerge between the third and fifth harmonics in HHG even in the simplest single-electron systems driven by coherent light, which is shown to be consistent across different values of the light-matter coupling $\lambda$ (see SI). This entanglement between harmonics is likely caused by intermode squeezing of the third and fifth harmonic, as previously suggested [21, 36].
We further examine slight variations of the Hamiltonian in Eq. (1) to test the generality of the trend observed in the averaged $R$ of Fig. 2. We will consider, on the one hand, different harmonics of the laser frequency emitted by the same atomic system and, on the other hand, different atomic models for the same third and fifth harmonics. Fig. 3 shows violation of Cauchy-Schwarz inequalities for the emitted 13th and 17th harmonics (top panel), and the 5th and the 17th harmonics (bottom panel); providing examples for both entanglement between above-threshold harmonics as well as between above- and below-threshold ones. These two scenarios exhibit a similar pattern to the below-threshold ones: tuning of the entanglement for specific ranges of the peak laser intensity. This suggests that in order to obtain entangled HHG photons a laser with moderate strong-field intensity is preferred (ranges $0.7\times 10^{14}~\mathrm{W/cm^{2}}$ to $1.5\times 10^{14}~\mathrm{W/cm^{2}}$), which should guide future experiments. In contrast, the $R$ parameter between the 5th and the 17th harmonic shows strong entanglement even for high laser power regimes. In Fig. 4 we also provide the calculated $R$ parameter for the third and fifth harmonics with modifications in the atomic model employed. The green curve (top-right panel) shows the results of placing an absorber at the quiver length $L_{Q}=E_{0}/\omega_{L}^{2}$, which effectively removes long trajectories from the simulation. Notably, the oscillations observed for the excursion-converged simulation (blue curve, top-left panel) disappear and the harmonics remain entangled even for higher intensity drivings. The red and cyan curves (bottom-left and right panels, respectively) correspond to a Gaussian potential reproducing the Neon ionization potential $V(x)=-1.17946\times e^{-0.347x^{2}}$ and a soft-Coulomb screened potential of the form $V(x)=-(1+e^{-x^{2}})/\sqrt{x^{2}+1.9396}$ reproducing the Helium ionization energy ($I_{p}=0.9037~\mathrm{a.u.}$), respectively. This investigation in fact reveals that entanglement in HHG can behave very differently depending on the type of system (e.g. if the electronic trajectories are localized or not, or if the Coulomb potential is screened or not), which hints towards potential transferability issues between theories and experiments performed in different samples. Ultimately though, the combination of the atomic system and laser intensity could be used also as sensitive tools for manipulating entanglement in high photon number high energy states.
Conclusions—We have developed an exact, fully quantum, theory of strong-field laser-matter interactions, without relying on semiclassical trajectory approximations. Our theory naturally captures the generation of entanglement between high-harmonic photons as observed experimentally. Using this approach to describe HHG driven by a coherent light source in a model 1D atom, we showed that true quantum correlations arise even when the system is driven by classical light, which is quantified by the violation of the intermode Cauchy-Schwarz inequality captured by the $R$ parameter. We further demonstrated that focal averaging over classical degrees of freedom in experimental settings is essential for recovering the correct entanglement signature, and in general for calculating quantum optical features of HHG such as correlations. This result should advance state-of-the-art theories used to analyze experiments. Our results show that the Cauchy–Schwarz inequalities are violated between the third and fifth harmonics only for a rather narrow range of peak laser intensities ($0.7\times 10^{14}~\mathrm{W/cm^{2}}$ to $1.5\times 10^{14}~\mathrm{W/cm^{2}}$), corresponding to Keldysh parameters  1-1.5. Moreover, the entanglement measures strongly oscillate with the driving laser power. This indicates that enhancement of entanglement in HHG should be possible via fine-tuning of optical laser power, which plays a dominant role. Additionally, we considered entanglement involving higher harmonics and found entanglement both among above-threshold harmonics and between below- and above-threshold harmonics, which we showed can have different characteristics. Lastly, by extending our analysis to different atomic systems, we showed that long electronic trajectories, and long-range behavior in general, can substantially alter the qualitative behavior of entangled HHG. Practically, this could lead to regimes of entangled harmonics in one system such as Ne, which would be non-entangled in another system, like He, or vice-versa. This result hints towards potential transferability issues in quantum HHG experiments and simulations performed in different targets, and should serve as a word of caution in analysis of measurements.
Looking forward, our findings provide a theoretical foundation for future research aiming to characterize the quantum features of HHG such as quantum state tomography of inter-harmonic squeezing [31, 52, 53], higher number of effective photon modes [54], superradiant emission of many-atoms and solid-state materials [55, 19], and electronic-correlations effects in the squeezing of the emitted harmonics [27]. Such effects could also be explored with QED based schemes like quantum electrodynamical density functional theory (QEDFT) [56, 57, 58]. These effects are also expected to advance emerging theories, and serve as a benchmark for testing quantum optical approximations.
Acknowledgements
This work was funded by the European Union under the ERC Synergy Grant UnMySt (HEU GA No. 101167294). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the European Research Council can be held responsible for them. This work was also supported by the Cluster of Excellence ‘Advanced Imaging of Matter’ (AIM), Grupos Consolidados (IT1453-22) and Deutsche Forschungsgemeinschaft (DFG) - SFB-925 - project 170620586. The Flatiron Institute is a division of the Simons Foundation. We acknowledge support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena. S.d.l.P. acknowledges support from International Max Planck Research School. We would like to thank Burak Gürlek, Michael Ruggenthaler, Frank Schlawin, David Theidel, Misha Ivanov, and Matan Even Tzur for interesting discussions. O.N. acknowledges support of the Young Faculty Award from the National Quantum Science and Technology program of Israel’s Council of Higher Education Planning and Budgeting Committee and support from The Technion Helen Diller Quantum Center.
Appendix A Discussion on Cauchy-Schwarz inequalities and non-classicality
In this section we explain the meaning of the violation of the Cauchy-Schwarz inequality referring to two photon modes [59, 39, 40]. A single-mode state of light can be fully described by its density operator represented using a Glauber-Sudarshan distribution:
[…]
where $P(\alpha)$ characterizes the state of light and $d^{2}\alpha\equiv d\Re{\left\{\alpha\right\}}d\Im{\left\{\alpha\right\}}$. Any expectation value that is normally-ordered can be evaluated in this representation substituting all annihilation (creation) operators $\hat{a}$ ($\hat{a}^{\dagger}$) with the complex phase-space variables $\alpha$ ($\alpha^{*}$), and then integrating using the appropiate weight given by $P(\alpha)$:
$\braket{\hat{a}^{\dagger n}\hat{a}^{m}}=\int P(\alpha)\alpha^{*n}\alpha^{m}~d^{2}\alpha.$ (9)
Any single-mode state is defined as classical if and only if $P(\alpha)$ behaves as a probability distribution; that is, is a non-negative, well-behaved function $P(\alpha)\geq 0$. For $P(\alpha)$ that are negative-valued or are more singular than the Dirac delta function, the system is considered quantum or non-classical.
In order to test the ‘quantum-ness of a state’ of a single-mode photon, we can use the Mandel-Q parameter defined as:
$Q=\braket{\hat{a}^{\dagger}\hat{a}}\left[g^{(2)}-1\right],$ (10)
where $g^{(2)}=\braket{\hat{a}^{\dagger 2}\hat{a}^{2}}/\braket{\hat{a}^{\dagger}\hat{a}}^{2}$ is the normalized second-order correlation function of the mode. For any classical state of light, in which $P(\alpha)$ is a well-behaved, non-negative probability distribution, the following Cauchy-Schwarz inequality is fulfilled:
$\braket{\hat{a}^{\dagger 2}\hat{a}^{2}}=\int P(\alpha)\left|\alpha\right|^{4}~d^{2}\alpha\geq\left[\int P(\alpha)\left|\alpha\right|^{2}~d^{2}\alpha\right]^{2}=\braket{\hat{a}^{\dagger}\hat{a}}^{2},$ (11)
corresponding to $g^{(2)}\geq 1$ and $Q\geq 0$. Therefore, any state of light in which $g^{(2)}<1$ or $Q<0$ is non-classical and is named antibunched. In the case of high-harmonic generation (HHG) driven a classical field, the emission appears to be always bunched ($g^{(2)}>1$, as shown in Fig. 2 [1, 12]). Therefore, it is impossible to characterize its non-classicality using the Mandel parameter.
For a two-mode system, non-classicality can instead be characterized through second-order intermodal correlation functions. This is precisely what the $R$ parameter measures. For a two-mode photon system, the density operator can be written as:
[…]
where $P_{AB}(\alpha,\beta)$ fully describes the two-mode system in the Glauber-Sudarshan representation. As in the one-mode case, classicality requires that $P_{AB}(\alpha,\beta)$ is a positive, well-behaved probability distribution. The single-mode Glauber-Sudarshan distributions can be recovered via partial tracing of the two-mode density operator: $P_{A}(\alpha)=\int P_{AB}(\alpha,\beta)d^{2}\beta$ and $P_{B}(\beta)=\int P_{AB}(\alpha,\beta)d^{2}\alpha$. If the two-mode state $P_{AB}(\alpha,\beta)$ is classical, then the following inequality holds:
$\braket{\hat{a}^{\dagger}\hat{a}\hat{b}^{\dagger}\hat{b}}=\int P_{AB}(\alpha,\beta)\left|\alpha\right|^{2}\left|\beta\right|^{2}~d^{2}\alpha d^{2}\beta\leq\sqrt{\left[\int P_{A}(\alpha)\left|\alpha\right|^{4}~d^{2}\alpha\right]\left[\int P_{B}(\beta)\left|\beta\right|^{4}~d^{2}\beta\right]}=\sqrt{\braket{\hat{a}^{\dagger 2}\hat{a}^{2}}\braket{\hat{b}^{\dagger 2}\hat{b}^{2}}}.$ (13)
Defining $R=\braket{\hat{a}^{\dagger}\hat{a}\hat{b}^{\dagger}\hat{b}}^{2}/\braket{\hat{a}^{\dagger 2}\hat{a}^{2}}\braket{\hat{b}^{\dagger 2}\hat{b}^{2}}$, any system with $R>1$ is non-classical (if we were talking about a many-boson system, then this would also entail multipartite entanglement). In our scenario, both partial Glauber-Sudarshan representations $P_{A}(\alpha)$ and $P_{B}(\beta)$ do not exhibit antibunching, so we cannot use them as tests for non-classicality in our theory. However, the combined distribution $P_{AB}(\alpha,\beta)$ does show non-classicality, as it violates the Cauchy-Schwarz distribution described in the main text. For Gaussian states or classical mixtures of Gaussian states (which is what perturbative analytical theories predict [36]) the condition $R>1$ is only satisfied if there is intermode squeezing or entanglement.
Appendix B Discussion on different values of the light-matter coupling
Fig. S1 presents the calculated $R$ parameter between the third and fifth harmonic modes for various values of the light-matter coupling strength, $\lambda=\sqrt{8\pi/L_{C}}$, where $L_{C}$ is the cavity length. The coupling strength is varied from $0.005~\mathrm{a.u.}$ to $0.02~\mathrm{a.u.}$. While the overall trend remains consistent across different coupling strengths, the degree of entanglement varies significantly: smaller light-matter couplings lead to higher maximum values of $R$. A more accurate characterization of this behavior would likely require methods capable of capturing a larger number of photon modes.
Figure S1: $R$ parameter for different values of the light-matter coupling $\lambda$ from $0.005$ to $0.02$. Left figure shows single-intensity simulations while right figure shows focal-averaged results.
These different light-matter couplings also correspond to different physical scenarios, as the mode volume is implicitly assumed in the evaluation of the observables (in this case, correlations between the modes). Specifically, each value of $\lambda$ reflects a different effective quantization length $L_{C}$ used in the calculation of the correlation function $g^{(2)}$.
In our simulations, we chose $\lambda=0.01~\mathrm{a.u.}$, as for a free-propagating mode with an envelope $f(t)$ defined in Eq. (2), the effective mode volume can be approximately determined by the time window of the envelope. Our finite pulse contains around 8 cycles of the fundamental frequency period $2\pi/\omega_{L}$, corresponding to an effective pulse length of approximately $L_{\mathrm{eff}}\approx 16\pi/\alpha\omega_{L}\sim 6.4~\mathrm{\mu m}$, where $\alpha$ is the fine-structure constant, which in atomic units corresponds to the speed of light. The cavity length chosen for our simulations corresponds to twice the effective distance the mode would need to propagate to be fully detected as a free-propagating wave $L_{C}=2L_{\mathrm{eff}}$, as we place the atom at the center of the cavity and we want the pulse to fully propagate through it.
Appendix C Technical aspects of the simulation
The envelope $f(t)$ presented in the Hamiltonian of Eq. (1) is the following:
$f(t)=\left[\Theta(t)-\Theta(t-t_{r})\right]\sin^{2}\left(\frac{\pi t}{2t_{r}}\right)+\Theta(t-t_{r})-\Theta(t-t_{d})+\left[\Theta(t-t_{d})-\Theta(t-t_{s})\right]\cos^{2}\left(\frac{\pi(t-t_{d})}{2t_{r}}\right),$ (14)
with $t_{r}=2.5\tau_{L}$, $t_{d}=5.5\tau_{L}$, $t_{s}=8\tau_{L}$, and $\tau_{L}=2\pi/\omega_{L}$.
The electron-photon wavefunction is represented in a three-dimensional real-space grid in which the coordinates x,y,zx,~y,~z represent the electron and the two photon modes, respectively. The computed photon wavefunction of the two modes $\phi(y,z;t)=\braket{y,z|\phi(t)}$ is recovered by projecting the of the electron-photon wavefunction $\Psi(x,y,z;t)=\braket{x,y,z|\Psi(t)}$ into the ground-state of the electronic system $\varphi^{*}_{g}(x)=\braket{g|x}$ at the end of the simulation $t_{F}>t_{s}$) is computed in the following way:
[…]
Appendix D Complete figures of photon observables used in the main text
Figure S2: Complete panel of photon observables for the 13th and 17th modes at the end of the simulation time for a Neon soft-Coulomb potential.
Figure S3: Complete panel of photon observables for the 5th and 17th modes at the end of the simulation time for a Neon soft-Coulomb potential.
Figure S4: Complete panel of photon observables for the 3rd and 5th modes at the end of the simulation time for a Neon soft-Coulomb potential and an absorber placed at the quiver length $L_{Q}=E_{0}/\omega_{L}^{2}$ that effectively removes long trajectories.
Figure S5: Complete panel of photon observables for the 3rd and 5th modes at the end of the simulation time for a Neon Gaussian potential $V(x)=-1.17946\times e^{-0.347x^{2}}$.
Figure S6: Complete panel of photon observables for the 3rd and 5th modes at the end of the simulation time for a Helium screened Gaussian potential $V(x)=-(1+e^{-x^{2}})/\sqrt{x^{2}+1.9396}$.